# Movable Antenna Enhanced Networked Full-Duplex Integrated Sensing and Communication System

### v2

@@ III-H Optimizing The RBS MAs’ Position
Therefore, the variable $\mathbf{t}^{1}_{m,n}$ optimization reduces to solving the following problem
[…]
$+\textrm{Re}\{(\mathbf{d}^{9}_{m,l,n})^{H}\mathbf{h}^{1}_{m,l,n}\}\big{)}+c_{5,1,m,n}$
$+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{(d^{11}_{j,m,n})^{\ast}\mathbf{g}_{r,m,j}[n]\}$
s.t. ${\sum}_{l=1}^{K_{u}}\big{(}(\mathbf{h}^{1}_{m,l,n})^{H}\mathbf{D}^{15}_{l,m,n,n}\mathbf{h}^{1}_{m,l,n}$ (76b)
$\!+\!\!\textrm{Re}\!\{\!(\mathbf{d}^{16}_{l,m,n})^{H}\!\mathbf{h}^{1}_{m,l,n}\}\!\big{)}\!\!+\!\!\!{\sum}_{j=1}^{K_{t}}\!\big{(}\!\textrm{Re}\{\!(d^{17}_{j,m,n})^{\ast}\!\mathbf{g}_{r,m,j}[n]\}\!\big{)}$
$-\textrm{Re}\{(d^{19}_{m,n})^{\ast}\mathbf{g}_{r,m,0}[n]\}+c_{5,2,m,n}\leq 0,$
$\|\mathbf{t}^{1}_{m,n}-\mathbf{t}^{1}_{m,i}\|_{2}\geq D_{t},\forall n,i\in\mathcal{N}_{t},n\neq i,$ (76c)
$\mathbf{t}^{1}_{m,n}\in\mathcal{C}.$ (76d)
It is obviously that the above problem is intractable. Next, we can use the MM framework to solve it. However, due to the structural similarity between problems (P9) and (P11), the convexity arguments utilized in (P9) are equally applicable to (P11), e.g., the convexification procedures for the objective function (76a) can be referenced in (61) and (64). Therefore, the specifics of these convexity operations, as detailed in Sec. III-G, are omitted here to avoid redundancy. And then, the optimization problem (P11) is rewritten as
[…]
s.t. $\tau_{2,m,n}\|\mathbf{t}^{1}_{m,n}\|^{2}_{2}\!+\!(\mathbf{d}^{21}_{m,n})^{T}\mathbf{t}^{1}_{m,n}\!+\!c_{5,4,m,n}\leq 0,$ (77b)
$\frac{(\mathbf{t}^{1}_{m,n,0}\!\!-\!\mathbf{t}^{1}_{m,i})^{T}}{\|\mathbf{t}^{1}_{m,n,0}\!\!-\!\mathbf{t}^{1}_{m,i}\|_{2}}\!(\mathbf{t}^{1}_{m,n}\!\!\!-\!\!\mathbf{t}^{1}_{m,i})\!\!\geq\!\!D_{t},\!\forall\!n,\!i\!\!\in\!\!\mathcal{N}_{r},n\!\!\neq\!\!i,$ (77c)
$\mathbf{t}^{1}_{m,n}\in\mathcal{C},$ (77d)
where $c_{5,3,m,n}$ and $c_{5,4,m,n}$ are constant terms, $\mathbf{t}^{1}_{m,n,0}$ is derived from the previous iteration, $\tau_{1,m,n}$ and $\tau_{2,m,n}$ are all positive real numbers and satisfy the condition in Lemma 1, and additional coefficients are omitted here for brevity (refer to Sec. III-G). The optimal solution of problem (P12) can be obtained by CVX.

### v3

@@ III-H Optimizing The RBS MAs’ Position
Therefore, the variable $\mathbf{t}^{1}$ optimization reduces to solving the following problem
$\textrm{(P11)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{t}^{1}}{\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$ (81a)
$-\textrm{Re}\{(\mathbf{{d}}^{2}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{(d^{3}_{j})^{\ast}\mathbf{g}_{r,j}\}+c_{5,1}$
s.t. ${\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)$ (81b)
$\!+\!{\sum}_{j=1}^{K_{t}}\!\big(\!\textrm{Re}\{\!(d^{5}_{j})^{\ast}\!{g}_{r,j}\}\!\big)-\textrm{Re}\{(d^{6})^{\ast}\mathbf{g}_{r,0}\}+c_{5,2}\leq 0,$
$\|\mathbf{t}^{1}-\mathbf{t}^{1}_{m,i}\|_{2}\geq D_{t},\forall n,i\in\mathcal{N}_{t},n\neq i,$ (81c)
$\mathbf{t}^{1}\in\mathcal{C}.$ (81d)
In the following, we will obtain the upper-bound of the objective function (81a) via convexify it. By leveraging the arguments presented in (62), a tight upper-bound of $(\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$ can be shown as follows
$(\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$ (82)
$\leq\lambda_{max}(\mathbf{D}^{1}_{l})(\|\mathbf{h}^{1}_{l}\|^{2}_{2}+\|\mathbf{h}^{1}_{l,0}\|^{2}_{2})-2\textrm{Re}\{\lambda_{max}(\mathbf{D}^{1}_{l})(\mathbf{h}^{1}_{l,0})^{H}\mathbf{h}^{1}_{l}\}$
$\!+\!2\textrm{Re }\{(\mathbf{h}^{1}_{l,0})^{H}\mathbf{D}^{1}_{l}(\mathbf{h}^{1}_{l}\!-\!\mathbf{h}^{1}_{l,0})\}+(\mathbf{h}^{1}_{l,0})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l,0}\!+\!\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$
$=\!\textrm{Re}\{(\mathbf{d}^{7}_{l})^{H}\mathbf{h}^{1}_{l}\}\!+\!c_{5,3,l}\!+\!\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\!\mathbf{h}^{1}_{l}\}\!=\!\textrm{Re}\{(\mathbf{d}^{8}_{l})^{H}\!\mathbf{h}^{1}_{l}\}\!\!+\!\!c_{5,3,l},$
where $\mathbf{t}^{1}_{0}$ is the feasible solution obtained in the last iteration, $\mathbf{h}^{1}_{l,0}\!\triangleq\!\mathbf{h}^{1}_{l}(\mathbf{t}^{1}_{0})$, $\|\mathbf{h}^{1}_{l}\|^{2}_{2}\!=\!\|\mathbf{h}^{1}_{l,0}\|^{2}_{2}\!=\!L^{t,1}_{m,l}$, $c_{5,3,l}\!\triangleq\!2\lambda_{max}(\mathbf{D}^{1}_{l})L^{t,1}_{m,l}-((\mathbf{h}^{1}_{l,0})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l,0})^{\ast}$, $\mathbf{d}^{7}_{l}\!\triangleq\!2((\mathbf{D}^{1}_{l})^{H}\mathbf{h}^{1}_{l,0}\!-\!\lambda_{max}(\mathbf{D}^{1}_{l})\mathbf{h}^{1}_{l,0})$ and $\mathbf{d}^{8}_{l}\!\triangleq\!\mathbf{d}^{7}_{l}\!+\!\mathbf{d}^{1}_{l}$.
Furthermore, by using the inequality (65) in Lemma 1, we can obtain a convex upper-bound of $\textrm{Re}\{(\mathbf{d}^{8}_{l})^{H}\mathbf{h}^{1}_{l}\}$ as follows
$\textrm{Re}\{(\mathbf{d}^{8}_{l})^{H}\mathbf{h}^{1}_{l}\}\triangleq h^{3}_{l}(\mathbf{t}^{1})$ (83)
$={\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{8}_{l}[i]|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{r,1}_{m,l,i}-\angle\mathbf{d}^{8}_{l}[i]\big)$
$\leq h^{3}_{l}(\mathbf{t}^{1}_{0})+\nabla^{T}h^{3}_{l}(\mathbf{t}^{1}_{0})(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})+{0.5\tau_{1,l}}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})^{T}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})$
$={0.5\tau_{1,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{9}_{l})^{T}\mathbf{t}^{1}+c_{5,4,l}$
where $c_{5,4,l}$ is a constant term, $\tau_{1,l}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{8}_{l}[i]|$, $\mathbf{d}^{9}_{l}\triangleq\nabla h^{3}_{l}(\mathbf{t}^{1}_{0})-\tau_{1,l}\mathbf{t}^{1}_{0}$, $\nabla h^{3}_{l}(\mathbf{t}^{1}_{0})$ is the gradient vector.
Furthermore, by leveraging the inequality (64), the lower-bound of $\textrm{Re}\{(\mathbf{d}^{2}_{l})^{H}\mathbf{h}^{1}_{l}\}$ is given as
$\textrm{Re}\{(\mathbf{d}^{2}_{l})^{H}\mathbf{h}^{1}_{l}\}\triangleq h^{4}_{l}(\mathbf{t}^{1})$ (84)
$={\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{2}_{l}[i]|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{r,1}_{m,l,i}-\angle\mathbf{d}^{2}_{l}[i]\big)$
$\geq h^{4}_{l}(\mathbf{t}^{1}_{0})+\nabla^{T}h^{4}_{l}(\mathbf{t}^{1}_{0})(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})-{0.5\bar{\tau}_{1,l}}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})^{T}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})$
$=-{0.5\bar{\tau}_{1,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{10}_{l})^{T}\mathbf{t}^{1}+c_{5,5,l},$
with $c_{5,5,l}$ being constant term, $\bar{\tau}_{1,l}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{2}_{l}[i]|$, $\mathbf{d}^{10}_{l}\triangleq\bar{\tau}_{1,l}\mathbf{t}^{1}_{0}+h^{4}_{l}(\mathbf{t}^{1}_{0})$, $\nabla h^{4}_{l}(\mathbf{t}^{1}_{0})$ is the gradient vector.
Next, the term $\textrm{Re}\{(d^{3}_{j})^{\ast}{g}_{r,j}\}$ can be convexified by leveraging (65) as follows
$\textrm{Re}\{(d^{3}_{j})^{\ast}{g}_{r,j}\}\triangleq g^{4}_{j}(\mathbf{t}^{1})$ (85)
$=|d^{3}_{j}||\mathbf{g}_{r,j}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{5}_{m,j}-\angle d^{3}_{j}\big)$
$\leq g^{4}_{j}(\mathbf{t}^{1}_{0})\!+\!\nabla^{T}g^{4}_{j}(\mathbf{t}^{1}_{0})(\mathbf{t}^{1}\!-\!\mathbf{t}^{1}_{0})\!+\!{0.5{\tau}_{2,j}}(\mathbf{t}^{1}\!-\!\mathbf{t}^{1}_{0})^{T}(\mathbf{t}^{1}\!-\!\mathbf{t}^{1}_{0})$
$={0.5\tau_{2,j}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{11}_{j})^{T}\mathbf{t}^{1}+c_{5,6,j},$
with $c_{5,6,j}$ being constant, $\tau_{2,j}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|d^{3}_{j}||\mathbf{g}_{r,j}|$, $\mathbf{d}^{11}_{j}\triangleq\nabla g^{4}_{j}(\mathbf{t}^{1}_{0})-\tau_{2,j}\mathbf{t}^{1}_{0}$, $\nabla g^{4}_{j}(\mathbf{t}^{1}_{0})$ is the gradient vector.
Based on the above transformation, the upper-bound of the objective function (81a) is given as
${\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$ (86)
$-\textrm{Re}\{(\mathbf{d}^{2}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{(d^{3}_{j})^{\ast}\mathbf{g}_{r,j}\}+c_{5,1}$
$\leq{\sum}_{l=1}^{K_{u}}\big({0.5\tau_{1,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{9}_{l})^{T}\mathbf{t}^{1}+c_{5,4,l}$
$-(-{0.5\bar{\tau}_{1,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{10}_{l})^{T}\mathbf{t}^{1}+\bar{c}_{5,5,l})\big)$
$+{\sum}_{j=0}^{K_{t}}(0.5{\tau_{2,j}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{11}_{j})^{T}\mathbf{t}^{1}+c_{5,6,j})+c_{5,1}$
$={\tau_{3}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{12})^{T}\mathbf{t}^{1}+c_{5,7},$
where $c_{5,7}$ is a constant term, $\tau_{3}\triangleq{\sum}_{l=1}^{K_{u}}\big({\tau_{1,l}}/{2}+{\bar{\tau}_{1,l}}/{2}\big)+{\sum}_{j=0}^{K_{t}}({\tau_{2,j}}/{2})$ and $\mathbf{d}^{12}\triangleq{\sum}_{l=1}^{K_{u}}(\mathbf{d}^{9}_{l}-\mathbf{d}^{10}_{l})+{\sum}_{j=0}^{K_{t}}\mathbf{d}^{11}_{j}$.
Next, we will construct the upper-bound of the sensing SINR constraint (81b). Firstly, by leveraging the arguments presented in (62) again, a tight upper-bound of $(\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}$, can be formulated as
$(\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}$ (87)
$\leq\lambda_{max}(\mathbf{D}^{2}_{l})(\|\mathbf{h}^{1}_{l}\|^{2}_{2}+\|\mathbf{h}^{1}_{l,0}\|^{2}_{2})-2\textrm{Re}\{\lambda_{max}(\mathbf{D}^{2}_{l})(\mathbf{h}^{1}_{l0})^{H}\mathbf{h}^{1}_{l}\}$
$+2\textrm{Re}\{(\mathbf{h}^{1}_{l,0})^{H}\mathbf{D}^{2}_{l}(\mathbf{h}^{1}_{l}\!-\!\mathbf{h}^{1}_{l,0})\}\!\!+\!\!(\mathbf{h}^{1}_{l,0})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l,0}\!\!+\!\!\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}$
$=\!\textrm{Re}\{(\mathbf{d}^{13}_{l})^{H}\!\mathbf{h}^{1}_{l}\}\!\!+\!\!c_{5,8,l}\!+\!\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\!\mathbf{h}^{1}_{l}\}\!\!=\!\textrm{Re}\{(\mathbf{d}^{14}_{l})^{H}\!\mathbf{h}^{1}_{l}\}\!\!+\!\!c_{5,8,l},$
where $c_{5,8,l}$ is constant, $\mathbf{d}^{13}_{l}\!\!\triangleq\!\!2((\mathbf{D}^{2}_{l})^{H}\mathbf{h}^{1}_{l,0}\!-\!\lambda_{max}(\mathbf{D}^{2}_{l})\mathbf{h}^{1}_{l,0})$ and $\mathbf{d}^{14}_{l}\!\triangleq\!\mathbf{d}^{13}_{l}\!+\!\mathbf{d}^{2}_{l}$.
Furthermore, by using the inequality (65) in Lemma 1, a convex upper-bound of $\textrm{Re}\{(\mathbf{d}^{14}_{l})^{H}\mathbf{h}^{1}_{l}\}$ is given as follows
$\textrm{Re}\{(\mathbf{d}^{14}_{l})^{H}\mathbf{h}^{1}_{l}\}\triangleq h^{5}_{l}(\mathbf{t}^{1})$ (88)
$={\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{14}_{l}[i]|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{r,1}_{m,l,i}-\angle\mathbf{d}^{14}_{l}[i]\big)$
$\leq h^{5}_{l}(\mathbf{t}^{1}_{0})+\nabla^{T}h^{5}_{l}(\mathbf{t}^{1}_{0})(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})+{0.5\tau_{4,l}}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})^{T}(\mathbf{t}^{1}-\mathbf{t}^{1}_{0})$
$={0.5\tau_{4,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{15}_{l})^{T}\mathbf{t}^{1}+c_{5,9,l},$
where $c_{5,9,l}$ is a constant term, $\tau_{4,l}=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{i=1}^{L^{t,1}_{m,l}}|\mathbf{d}^{14}_{l}[i]|$, $\mathbf{d}^{15}_{l}\triangleq\nabla h^{5}_{l}(\mathbf{t}^{1}_{0})-\tau_{4,l}\mathbf{t}^{1}_{0}$, $\nabla h^{5}_{l}(\mathbf{t}^{1}_{0})$ is the gradient vector.
Similarly, the term $\textrm{Re}\{(d^{5}_{j})^{\ast}{g}_{r,j}\}$ can be convexified by leveraging (65) as follows
$\textrm{Re}\{(d^{5}_{j})^{\ast}{g}_{r,j}\}\triangleq g^{5}_{j}(\mathbf{t}^{1})$ (89)
$=|d^{5}_{j}||\mathbf{g}_{r,j}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{5}_{m,j}-\angle d^{5}_{j}\big)$
$\leq{0.5\tau_{5,j}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{16}_{j})^{T}\mathbf{t}^{1}+c_{5,10,j},$
with $c_{5,10,j}$ being constants, $\tau_{5,j}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|d^{5}_{j}||{g}_{r,j}|$, $\mathbf{d}^{16}_{j}\triangleq\nabla g^{5}_{j}(\mathbf{t}^{1}_{0})-\tau_{5,j}\mathbf{t}^{1}_{0}$, $\nabla g^{5}_{j}(\mathbf{t}^{1}_{0})$ is the gradient vector.
And then, a convex lower bound for $\text{Re}\{(d^{6})^{\ast}{g}_{r,0}\}$ by applying (64) in Lemma 1 is given as follows
$\text{Re}\{(d^{6})^{\ast}{g}_{r,0}(\mathbf{t}^{1})\}\triangleq g^{6}(\mathbf{t}^{1})$ (90)
$=|d^{6}||\mathbf{g}_{r,0}|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{1})^{T}\mathbf{a}^{5}_{m,0}-\angle d^{6}\big)$
$\geq-{0.5\tau_{6}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{17})^{T}\mathbf{t}^{1}_{m,n}+c_{5,11},$
where $c_{5,11}$ is constant, $\tau_{6}\triangleq\frac{8\pi^{2}}{\lambda^{2}}|d^{6}||\mathbf{g}_{r,0}|$, $\mathbf{d}^{17}\triangleq\nabla g^{6}(\mathbf{t}^{1})+\tau_{6}\mathbf{t}^{1}_{0}$, $\nabla g^{6}(\mathbf{t}^{0})$ is the the gradient vector.
Therefore, a convex upper-bound of the sensing SINR constraint (81b) can be expressed as
${\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)$ (91)
$+{\sum}_{j=1}^{K_{t}}\big(\textrm{Re}\{(d^{5}_{j})^{\ast}{g}_{r,j}\}\big)-\textrm{Re}\{(d^{6})^{\ast}\mathbf{g}_{r,0}\}+c_{5,2}$
$\leq{\sum}_{l=1}^{K_{u}}\big({0.5\tau_{4,l}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{15}_{l})^{T}\mathbf{t}^{1}+c_{5,9,l}+c_{5,8,l}\big)$
$+{\sum}_{j=1}^{K_{t}}({0.5\tau_{5,j}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{16}_{j})^{T}\mathbf{t}^{1}+c_{5,10,j})$
$-(-{0.5\tau_{6}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{17})^{T}\mathbf{t}^{1}_{m,n}+c_{5,11})+c_{5,2},$
$=\tau_{7}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{18})^{T}\mathbf{t}^{1}+c_{5,12},$
where $c_{5,12}$ is constant, $\tau_{7}\triangleq{\sum}_{l=1}^{K_{u}}({\tau_{4,l}}/{2})+{\sum}_{j=1}^{K_{t}}({\tau_{5,j}}/{2})-(-{\tau_{6}}/{2})$, $\mathbf{d}^{18}\triangleq{\sum}_{l=1}^{K_{u}}\mathbf{d}^{15}_{l}+{\sum}_{j=1}^{K_{t}}\mathbf{d}^{16}_{j}-(\mathbf{d}^{17})$.
Inspired by the MM framework, the constraint (81c) can be linearized as follows
$\|\mathbf{t}^{1}-\mathbf{t}^{1}_{m,i}\|_{2}\geq\frac{(\mathbf{t}^{1}_{0}-\mathbf{t}^{1}_{m,i})^{T}}{\|\mathbf{t}^{1}_{0}-\mathbf{t}^{1}_{m,i}\|_{2}}(\mathbf{t}^{1}-\mathbf{t}^{1}_{m,i}).$ (92)
Therefore, the optimization problem (P11) is rewritten as
$\textrm{(P12)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{t}^{1}}{\tau_{3}}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{12})^{T}\mathbf{t}^{1}+c_{5,7}$ (93a)
s.t. $\tau_{7}\|\mathbf{t}^{1}\|^{2}_{2}+(\mathbf{d}^{18})^{T}\mathbf{t}^{1}+c_{5,12}\leq 0,$ (93b)
$\frac{(\mathbf{t}^{1}_{0}-\mathbf{t}^{1}_{m,i})^{T}}{\|\mathbf{t}^{1}_{0}-\mathbf{t}^{1}_{m,i}\|_{2}}(\mathbf{t}^{1}-\mathbf{t}^{1}_{m,i})\geq\!\!D_{t},\!\forall\!n,\!i\!\in\!\mathcal{N}_{r},n\!\neq\!i,$ (93c)
$\mathbf{t}^{1}\in\mathcal{C}.$ (93d)
Obviously, the problem (P12) is convex and its optimal solution can be obtained by CVX.